### 10 Apr 2026 v1

# From Distance to Angle: One-Shot Detection Under Additive White Cauchy Noise

Abstract
We study one-shot detection under additive white Cauchy noise (AWCN) using finite constellations, with emphasis on the geometric mechanisms governing symbol-level reliability. Under isotropic Cauchy noise, the maximum-likelihood rule induces the same Euclidean Voronoi decision regions as in the Gaussian case, so the distinction lies not in the decision geometry itself but in how probability mass is distributed over these fixed regions. In the small-noise regime, we derive a reciprocal distance-spectrum upper bound for the symbol error probability, showing that reliability retains a longer-range dependence on the global constellation geometry than under additive white Gaussian noise. In the large-noise regime, we prove that the correct-decision probability converges to a limit determined solely by the angular measure of the associated Voronoi recession cone. These results formalize a regime-dependent transition from distance-based to angle-based reliability descriptors under heavy-tailed noise. The theory is further illustrated through an asymmetric four-point example exhibiting geometric collapse and a standard 4QAM sanity check.
I Introduction
Additive white Gaussian noise (AWGN) has long served as the canonical model for detection and constellation design, under which reliability is largely governed by minimum-distance geometry. This viewpoint underlies much of modern digital communication theory [9], particularly in the small-noise regime. Recent studies, however, have identified communication and sensing scenarios in which noise exhibits heavy-tailed, non-Gaussian behavior, rendering Gaussian-based design principles suboptimal [10, 6]. Such heavy-tailed noise models have also attracted recent interest in information-theoretic studies [1, 15, 8].
Specifically, the Cauchy model has emerged as a prototypical member of the $\alpha$-stable family [6], offering a mathematically rigorous yet tractable departure from the Gaussian assumption. While its infinite variance and polynomial tail decay complicate traditional Law-of-Large-Numbers-based analysis [15], the model is far from a mere theoretical abstraction. It arises naturally in practical systems, including massive MIMO [4] and diffusion-based molecular communication [2, 5], where nearest-neighbor geometry no longer suffices to characterize reliability. This necessitates identifying the alternative geometric descriptors under heavy-tailed noise, an area partially explored in [11, 13, 12].
On the other hand, recent findings [8] reveal that channels subject to AWCN can sustain positive information rates under a second-moment input (energy) constraint. By positioning AWCN as a meaningful benchmark directly comparable to AWGN under identical energy normalization, these results motivate a timely re-examination of finite-constellation detection and design [14, 9] in the presence of Cauchy noise.
To address this need, this paper formally develops a geometric asymptotic theory for one-shot detection over AWCN channels. Our focus is not on modifying the decision rule itself, which remains Euclidean under isotropic Cauchy noise, but on identifying the geometric descriptors that govern reliability across different noise regimes. In the small-noise regime, we derive a reciprocal distance-spectrum upper bound showing that one-shot reliability retains a longer-range sensitivity to the global constellation geometry than in the Gaussian case. In the large-noise regime, we prove that bounded distance information disappears asymptotically and that the correct-decision probability converges to a limit determined solely by the angular measure of the associated Voronoi recession cone (see [7]). Taken together, these results provide a theorem-based account of how one-shot reliability shifts from distance-dominated behavior to angle-dominated behavior under heavy-tailed noise.
The remainder of this paper is organized as follows. Section II introduces the system model and problem statement. Section III presents the small-noise distance-spectrum bound. Section IV establishes the large-noise angular convergence theorem and the associated geometric-collapse criterion. The main results are further illustrated through an asymmetric four-point worked example in Appendix A and a standard 44QAM sanity check in Appendix B. Conclusions are drawn in Section V.
II System Model and Problem Statement
II-A Vector Cauchy Channels
We focus on one-shot detection (also referred to as symbol-by-symbol detection), where the receiver performs optimal maximum-likelihood (ML) decoding based solely on the current observation vector. This setting isolates the symbol-level effect of heavy-tailed noise on decision geometry and reliability, without invoking coding across multiple channel uses.
We consider a $d$-dimensional discrete-time vector channel model defined as
$\bm{y}=\bm{x}+\bm{n},$ (1)
where $\bm{x}\in\mathcal{C}\subset\mathbb{R}^{d}$ is the transmitted signal vector from a finite constellation subject to an average energy constraint $\mathbb{E}[\|\bm{x}\|^{2}]\leq P$. The additive noise $\bm{n}\in\mathbb{R}^{d}$ is modeled as an isotropic $d$-dimensional Cauchy random vector with probability density function (PDF) (see [10, 15]),
$f_{\bm{N}}(\bm{n})=c_{d}\,\gamma^{-d}\left(1+\frac{\|\bm{n}\|^{2}}{\gamma^{2}}\right)^{-\frac{d+1}{2}},\quad c_{d}=\frac{\Gamma\!\left(\frac{d+1}{2}\right)}{\pi^{\frac{d+1}{2}}},$ (2)
where $\gamma>0$ is the scale parameter controlling the distribution’s width. Unlike Gaussian noise, the Cauchy distribution lacks finite moments; thus, standard signal-to-noise ratio (SNR) is technically undefined. We instead use the normalized proxy $P/\gamma^{2}$ to characterize system performance (see [8]).
Under this spherically symmetric model, the ML decoding rule minimizes the negative log-likelihood, which simplifies gracefully to the standard Euclidean distance metric. Namely,
$\hat{\bm{x}}=\operatorname*{arg\,min}_{\bm{c}\in\mathcal{C}}\log\left(1+\frac{\|\bm{y}-\bm{c}\|^{2}}{\gamma^{2}}\right)\equiv\operatorname*{arg\,min}_{\bm{c}\in\mathcal{C}}\|\bm{y}-\bm{c}\|.$ (3)
Because the isotropic Cauchy density is monotonically decreasing with the Euclidean norm, the exact decision boundaries induced by (3) correspond to Euclidean Voronoi cells (see [7]), entirely independent of $\gamma$. Under equiprobable signaling, the average symbol error probability is formally given by
$P_{e}=1-\frac{1}{M}\sum_{\bm{x}\in\mathcal{C}}\int_{V(\bm{x})}f_{\bm{N}}(\bm{y}-\bm{x})\mathrm{d}\bm{y}.$ (4)
For the rest of this paper, we use uppercase letters such as $X,Y,N$ for random variables and lowercase letters such as $x,y,n$ for their realizations.
II-B The Combining Gain Paradox
In traditional AWGN vector channels, linear aggregation strategies (e.g., maximum ratio combining) yield a genuine averaging effect, so reliability typically improves with the number of spatial dimensions or channel paths [9]. Under AWCN, this self-averaging mechanism breaks down: the Cauchy law has neither finite mean nor finite variance, so the usual Law-of-Large-Numbers (LLN) intuition no longer applies. More fundamentally, because the Cauchy distribution is $1$-stable (see [10, 6]), any linear combination of independent Cauchy variables remains Cauchy. Consequently, linear combining does not concentrate the effective noise distribution, but merely produces another scalar Cauchy projection, thereby losing the classical multidimensional combining gain.
Because the exact $d$-dimensional Cauchy integral in (4) lacks a closed-form solution over arbitrary Voronoi polytopes, resolving this paradox requires moving beyond linear operations and standard distance formulations. The specific objective of this paper is to geometrically reinterpret the error integral (4). By establishing how symbol reliability transitions from distance-based to angle-based mechanisms, we aim to identify the geometric descriptors that govern one-shot detection under AWCN and that may inform future geometry-aware receiver design.
III Small-Noise Regime: Distance Spectrum Bounds
In this section, we formally characterize the one-shot reliability in the small-noise regime.
III-A Pairwise Comparison Term
Before deriving the small-noise bound, we isolate the basic pairwise comparison event that enters the union-bound argument. Fix two distinct constellation points $x_{i},x_{j}\in\mathcal{C}$, and define
$d_{ij}:=\|x_{i}-x_{j}\|,\qquad u_{ij}:=\frac{x_{j}-x_{i}}{\|x_{j}-x_{i}\|}.$ (5)
Assume that $x_{i}$ is transmitted. Since the isotropic Cauchy law is absolutely continuous, tie events occur with probability zero and do not affect the ML analysis. Under ML decoding, the event that $x_{j}$ is at least as favorable as $x_{i}$ is
$\|Y-x_{j}\|\leq\|Y-x_{i}\|.$ (6)
Since $Y=x_{i}+N$, this is equivalent to
$\|N-(x_{j}-x_{i})\|\leq\|N\|,$ (7)
or, after expanding the squared norms,
$u_{ij}^{\mathsf{T}}N\geq\frac{d_{ij}}{2}.$ (8)
Because the noise vector $N$ is isotropic Cauchy, every one-dimensional projection onto a unit direction remains a scalar Cauchy random variable with the same scale parameter $\gamma$. Therefore,
$\mathbb{P}\!\left(\|Y-x_{j}\|\leq\|Y-x_{i}\|\,\middle|\,X=x_{i}\right)$ $=\mathbb{P}\!\left(u_{ij}^{\mathsf{T}}N\geq\frac{d_{ij}}{2}\right)$ (9)
$=\frac{1}{2}-\frac{1}{\pi}\arctan\!\left(\frac{d_{ij}}{2\gamma}\right).$
Equation (9) shows that each pairwise comparison term depends only on the Euclidean distance $d_{ij}$. This quantity will serve as the basic building block in the distance-spectrum bound of Theorem 1.
III-B Main Result: Small-Noise Bound and Asymptotics
We define the distance spectrum of $\mathcal{C}$ as the multiset $\mathcal{S}(\mathcal{C})=\{d_{ij}:1\leq i<j\leq M\}$. The following theorem establishes the asymptotic behavior of the error probability as the noise scale vanishes.
Theorem 1 (Small-Noise Distance-Spectrum Bound).
Let $\mathcal{C}=\{x_{1},\dots,x_{M}\}\subset\mathbb{R}^{d}$ be a finite constellation, and consider the isotropic $d$-dimensional Cauchy channel
$Y=X+N,$ (10)
with maximum-likelihood decoding as in (3). For a fixed transmitted symbol $x_{i}\in\mathcal{C}$, define
$d_{ij}:=\|x_{i}-x_{j}\|,\qquad j\neq i.$ (11)
Then, for every $\gamma>0$, the conditional symbol error probability satisfies
$P_{e}(x_{i};\gamma)\leq\sum_{j\neq i}\left[\frac{1}{2}-\frac{1}{\pi}\arctan\!\left(\frac{d_{ij}}{2\gamma}\right)\right].$ (12)
Consequently, as $\gamma\to 0$,
$P_{e}(x_{i};\gamma)\leq\frac{2\gamma}{\pi}\sum_{j\neq i}\frac{1}{d_{ij}}+O(\gamma^{3}).$ (13)
If the symbols are equiprobable, then the average symbol error probability satisfies
$P_{e}(\gamma)\leq\frac{4\gamma}{M\pi}\sum_{1\leq i<j\leq M}\frac{1}{d_{ij}}+O(\gamma^{3}).$ (14)
Hence, in the small-noise regime, the one-shot reliability is controlled, at leading order, by the reciprocal distance spectrum of the constellation.
Proof.
Fix $x_{i}\in\mathcal{C}$ and assume that $x_{i}$ is transmitted. Let
$E_{i}:=\{\hat{X}\neq x_{i}\}$ (15)
denote the decoding error event under ML decoding.
For each competitor $x_{j}\in\mathcal{C}\setminus\{x_{i}\}$, define
$H_{ij}:=\Bigl\{n\in\mathbb{R}^{d}:\|x_{i}+n-x_{j}\|\leq\|n\|\Bigr\}.$ (16)
The event $H_{ij}$ means that, after adding noise $n$, the point $x_{j}$ is at least as close to the received vector as $x_{i}$. Therefore,
$E_{i}\subseteq\bigcup_{j\neq i}H_{ij},$ (17)
and hence, by the union bound,
$P_{e}(x_{i};\gamma)\leq\sum_{j\neq i}\mathbb{P}(H_{ij}).$ (18)
We now simplify each event $H_{ij}$. Expanding the squared norms gives
$\|x_{i}+n-x_{j}\|^{2}\leq\|n\|^{2}$ (19)
$\Longleftrightarrow\quad\|n-(x_{j}-x_{i})\|^{2}\leq\|n\|^{2},$
which is equivalent to
$2(x_{j}-x_{i})^{\mathsf{T}}n+\|x_{j}-x_{i}\|^{2}\leq 0.$ (20)
Thus,
$H_{ij}=\left\{(x_{j}-x_{i})^{\mathsf{T}}N\geq\frac{\|x_{j}-x_{i}\|^{2}}{2}\right\}.$ (21)
Let
$u_{ij}:=\frac{x_{j}-x_{i}}{\|x_{j}-x_{i}\|}=\frac{x_{j}-x_{i}}{d_{ij}}.$ (22)
Then
$H_{ij}=\left\{u_{ij}^{\mathsf{T}}N\geq\frac{d_{ij}}{2}\right\}.$ (23)
Because $N$ is an isotropic multivariate Cauchy vector with scale parameter $\gamma$, every one-dimensional projection $u^{\mathsf{T}}N$ onto a unit vector $u$ is a one-dimensional Cauchy random variable with location 0 and scale $\gamma$, i.e.,
$u^{\mathsf{T}}N\sim\mathrm{Cauchy}(0,\gamma),$ (24)
whose density is
$f_{Z}(z)=\frac{1}{\pi\gamma}\frac{1}{1+(z/\gamma)^{2}},\qquad z\in\mathbb{R}.$ (25)
Therefore,
$\mathbb{P}(H_{ij})$ $=\mathbb{P}\!\left(u_{ij}^{\mathsf{T}}N\geq\frac{d_{ij}}{2}\right)$ (26)
$=\int_{d_{ij}/2}^{\infty}\frac{1}{\pi\gamma}\frac{1}{1+(z/\gamma)^{2}}\,\mathrm{d}z.$
Evaluating the integral yields
$\mathbb{P}(H_{ij})=\frac{1}{2}-\frac{1}{\pi}\arctan\!\left(\frac{d_{ij}}{2\gamma}\right).$ (27)
Substituting this into (18) proves (12).
To obtain the small-noise expansion, use the standard asymptotic formula [3]:
$\arctan t=\frac{\pi}{2}-\frac{1}{t}+\frac{1}{3t^{3}}+O(t^{-5}),\qquad t\to\infty.$ (28)
Setting $t=d_{ij}/(2\gamma)$ gives, for each fixed $i\neq j$,
$\frac{1}{2}-\frac{1}{\pi}\arctan\!\left(\frac{d_{ij}}{2\gamma}\right)=\frac{2\gamma}{\pi d_{ij}}+O(\gamma^{3}),\qquad\gamma\to 0.$ (29)
Since the constellation is finite, summing over $j\neq i$ yields
$P_{e}(x_{i};\gamma)\leq\frac{2\gamma}{\pi}\sum_{j\neq i}\frac{1}{d_{ij}}+O(\gamma^{3}),$ (30)
which proves (13).
Finally, if the input symbols are equiprobable, then
$P_{e}(\gamma)=\frac{1}{M}\sum_{i=1}^{M}P_{e}(x_{i};\gamma).$ (31)
Averaging the previous bound over $i$ gives
$P_{e}(\gamma)\leq\frac{2\gamma}{M\pi}\sum_{i=1}^{M}\sum_{j\neq i}\frac{1}{d_{ij}}+O(\gamma^{3}).$ (32)
Using the symmetry $d_{ij}=d_{ji}$ and double counting,
$\sum_{i=1}^{M}\sum_{j\neq i}\frac{1}{d_{ij}}=2\sum_{1\leq i<j\leq M}\frac{1}{d_{ij}},$ (33)
which proves (14). ∎
Remark.
Theorem 1 clarifies an important contrast between Gaussian and Cauchy detection. Under AWGN, pairwise error terms decay exponentially with distance, so the nearest-neighbor structure typically dominates the small-noise behavior. Under AWCN, however, the pairwise comparison term in (9) satisfies
$\frac{1}{2}-\frac{1}{\pi}\arctan\!\left(\frac{d_{ij}}{2\gamma}\right)=\frac{2\gamma}{\pi d_{ij}}+O(\gamma^{3}),\qquad\gamma\to 0,$ (34)
which decays only algebraically in the distance $d_{ij}$. As a result, farther constellation points are suppressed more weakly than in the Gaussian case, and the small-noise error bound naturally depends on the full reciprocal distance spectrum rather than solely on the minimum distance. In this sense, Theorem 1 should be interpreted as a geometric upper bound showing that, even in the high-reliability regime, Cauchy noise preserves a longer-range sensitivity to the global constellation geometry.
IV Large-Noise Regime: Angular Voronoi Convergence
As $\gamma$ increases, the distance-based description breaks down. In this regime, the received vector is pushed far away from the transmitted symbol, so the bounded offsets among constellation points become negligible relative to the overall noise magnitude. We show that the decoding problem then loses sensitivity to finite Euclidean distances and retains only the asymptotic angular geometry of the Voronoi cells. The appropriate descriptor is therefore the recession cone of each Voronoi region and its associated spherical angular patch.
IV-A Voronoi Cones and Angular Patches
For each $x\in\mathcal{C}$, the associated Voronoi cell is
$V(x)=\{y\in\mathbb{R}^{d}:\ \|y-x\|\leq\|y-c\|,\ \forall c\in\mathcal{C}\}.$ (35)
Its recession cone is defined by
$\mathsf{K}(x)=\bigcap_{c\in\mathcal{C}\setminus\{x\}}\{u\in\mathbb{R}^{d}:\ u^{\mathsf{T}}(c-x)\leq 0\}.$ (36)
The spherical cross-section
$A_{x}:=\mathsf{K}(x)\cap\mathbb{S}^{d-1}$ (37)
is called the Voronoi angular patch of $x$, where $\mathbb{S}^{d-1}$ denotes the unit sphere in $\mathbb{R}^{d}$.
The geometric meaning of (36) is transparent after translating the Voronoi cell to the origin. Indeed,
$V(x)-x=\bigcap_{c\in\mathcal{C}\setminus\{x\}}\left\{z\in\mathbb{R}^{d}:\ z^{\mathsf{T}}(c-x)\leq\frac{\|c-x\|^{2}}{2}\right\}.$ (38)
Hence, after scaling by $\gamma$,
$\frac{V(x)-x}{\gamma}=\bigcap_{c\in\mathcal{C}\setminus\{x\}}\left\{u\in\mathbb{R}^{d}:\ u^{\mathsf{T}}(c-x)\leq\frac{\|c-x\|^{2}}{2\gamma}\right\},$ (39)
which suggests the recession-cone limit (36) as $\gamma\to\infty$. Thus, the Voronoi cone $\mathsf{K}(x)$ is precisely the large-noise limit of the translated-and-rescaled decision region.
IV-B Main Result: Large-Noise Convergence
The following theorem shows that, in the infinite-noise limit, the correct-decision probability depends only on the angular measure of the recession cone.
Theorem 2 (Large-Noise Convergence).
Let $\mathcal{C}\subset\mathbb{R}^{d}$ be a finite constellation, and let
$P_{c}(x;\gamma):=\mathbb{P}(\hat{X}=x\,|\,X=x)$ (40)
denote the conditional correct-decision probability under the isotropic $d$-dimensional Cauchy channel with scale parameter $\gamma$. For each $x\in\mathcal{C}$, let $\mathsf{K}(x)$ and $A_{x}$ be defined by (36) and (37), and let $\sigma_{d-1}$ denote the usual surface measure on $\mathbb{S}^{d-1}$. Then
$\lim_{\gamma\to\infty}P_{c}(x;\gamma)=\frac{\sigma_{d-1}(A_{x})}{\sigma_{d-1}(\mathbb{S}^{d-1})}.$ (41)
Equivalently,
$\lim_{\gamma\to\infty}P_{e}(x;\gamma)=1-\frac{\sigma_{d-1}(A_{x})}{\sigma_{d-1}(\mathbb{S}^{d-1})},$ (42)
where $P_{e}(x;\gamma)=1-P_{c}(x;\gamma)$.
If the constellation symbols are equiprobable, define
$P_{c}(\gamma):=\frac{1}{M}\sum_{x\in\mathcal{C}}P_{c}(x;\gamma).$ (43)
Then
$\lim_{\gamma\to\infty}P_{c}(\gamma)=\frac{1}{M}\sum_{x\in\mathcal{C}}\frac{\sigma_{d-1}(A_{x})}{\sigma_{d-1}(\mathbb{S}^{d-1})}.$ (44)
Proof.
Fix $x\in\mathcal{C}$. By definition of ML decoding,
$P_{c}(x;\gamma)=\mathbb{P}(Y\in V(x)\mid X=x)=\mathbb{P}(N\in V(x)-x).$ (45)
Since the noise density is
$f_{N}(n)=c_{d}\,\gamma^{-d}\left(1+\frac{\|n\|^{2}}{\gamma^{2}}\right)^{-\frac{d+1}{2}},$ (46)
we obtain
$P_{c}(x;\gamma)=\int_{V(x)-x}c_{d}\,\gamma^{-d}\left(1+\frac{\|n\|^{2}}{\gamma^{2}}\right)^{-\frac{d+1}{2}}\,\mathrm{d}n.$ (47)
Applying the change of variables $n=\gamma u$ yields
$P_{c}(x;\gamma)=\int_{(V(x)-x)/\gamma}c_{d}\,(1+\|u\|^{2})^{-\frac{d+1}{2}}\,\mathrm{d}u.$ (48)
We now show that the indicator of the scaled region converges almost everywhere to the indicator of the recession cone. From (39),
$\frac{V(x)-x}{\gamma}=\bigcap_{c\in\mathcal{C}\setminus\{x\}}\left\{u:\ u^{\mathsf{T}}(c-x)\leq\frac{\|c-x\|^{2}}{2\gamma}\right\}.$ (49)
Let
$B_{x}:=\bigcup_{c\in\mathcal{C}\setminus\{x\}}\{u\in\mathbb{R}^{d}:\ u^{\mathsf{T}}(c-x)=0\}.$ (50)
Since $\mathcal{C}$ is finite, $B_{x}$ is a finite union of hyperplanes and therefore has Lebesgue measure zero.
Now fix $u\notin B_{x}$. Then for every $c\neq x$, the quantity $u^{\mathsf{T}}(c-x)$ is nonzero. If $u\in\mathsf{K}(x)$, then all these inner products are strictly negative, so for sufficiently large $\gamma$ they satisfy
$u^{\mathsf{T}}(c-x)\leq\frac{\|c-x\|^{2}}{2\gamma},\qquad\forall c\neq x,$ (51)
and hence $u\in(V(x)-x)/\gamma$. Conversely, if $u\notin\mathsf{K}(x)$, then there exists some $c\neq x$ such that $u^{\mathsf{T}}(c-x)>0$, and for sufficiently large $\gamma$,
$u^{\mathsf{T}}(c-x)>\frac{\|c-x\|^{2}}{2\gamma},$ (52)
so $u\notin(V(x)-x)/\gamma$. Therefore,
$\mathbf{1}_{(V(x)-x)/\gamma}(u)\to\mathbf{1}_{\mathsf{K}(x)}(u)\qquad\text{for a.e. }u\in\mathbb{R}^{d}.$ (53)
Since
$0\leq\mathbf{1}_{(V(x)-x)/\gamma}(u)\,c_{d}(1+\|u\|^{2})^{-\frac{d+1}{2}}\leq c_{d}(1+\|u\|^{2})^{-\frac{d+1}{2}},$ (54)
and the dominating function on the right is integrable over $\mathbb{R}^{d}$, the dominated convergence theorem applied to (48) and (53) gives
$\lim_{\gamma\to\infty}P_{c}(x;\gamma)=\int_{\mathsf{K}(x)}c_{d}(1+\|u\|^{2})^{-\frac{d+1}{2}}\,\mathrm{d}u.$ (55)
It remains to evaluate the right-hand side in angular form. Because the density is radial, we use polar coordinates $u=r\theta$, where $r\in[0,\infty)$ and $\theta\in\mathbb{S}^{d-1}$. Since $\mathsf{K}(x)$ is a cone, its polar decomposition is exactly
$\mathsf{K}(x)=\{r\theta:\ r\geq 0,\ \theta\in A_{x}\}.$ (56)
Hence
$\int_{\mathsf{K}(x)}c_{d}(1+\|u\|^{2})^{-\frac{d+1}{2}}\mathrm{d}u$ (57)
$=\int_{A_{x}}\!\int_{0}^{\infty}c_{d}(1+r^{2})^{-\frac{d+1}{2}}r^{d-1}\mathrm{d}r\mathrm{d}\sigma_{d-1}(\theta)$
$=\sigma_{d-1}(A_{x})\underbrace{\int_{0}^{\infty}c_{d}(1+r^{2})^{-\frac{d+1}{2}}r^{d-1}\mathrm{d}r}_{=:I_{d}}.$
On the other hand, integrating the same radial density over the entire space $\mathbb{R}^{d}$ gives
$1=\int_{\mathbb{R}^{d}}c_{d}(1+\|u\|^{2})^{-\frac{d+1}{2}}\,\mathrm{d}u=\sigma_{d-1}(\mathbb{S}^{d-1})\,I_{d}.$ (58)
Thus
$I_{d}=\frac{1}{\sigma_{d-1}(\mathbb{S}^{d-1})}.$ (59)
Substituting this into (57) and then into (55) yields
$\lim_{\gamma\to\infty}P_{c}(x;\gamma)=\frac{\sigma_{d-1}(A_{x})}{\sigma_{d-1}(\mathbb{S}^{d-1})},$ (60)
which proves (41). Equation (42) follows from $P_{e}(x;\gamma)=1-P_{c}(x;\gamma)$, and (44) follows by averaging over equiprobable symbols. ∎
An immediate and striking consequence of Theorem 2 is the identification of constellation points that become entirely unreliable.
Corollary 1 (Condition for Geometric Collapse).
Let $x\in\mathcal{C}$. If the angular patch $A_{x}$ has zero surface measure, i.e.,
$\sigma_{d-1}(A_{x})=0,$ (61)
then
$\lim_{\gamma\to\infty}P_{c}(x;\gamma)=0,\qquad\lim_{\gamma\to\infty}P_{e}(x;\gamma)=1.$ (62)
In particular, if $x$ lies in the interior of $\operatorname{conv}(\mathcal{C})$, then $A_{x}=\varnothing$, and therefore geometric collapse occurs.
Proof.
If $\sigma_{d-1}(A_{x})=0$, then (41) in Theorem 2 immediately gives
$\lim_{\gamma\to\infty}P_{c}(x;\gamma)=\frac{\sigma_{d-1}(A_{x})}{\sigma_{d-1}(\mathbb{S}^{d-1})}=0,$ (63)
which proves (62).
For the interior-point claim, suppose that $x\in\operatorname{int}(\operatorname{conv}(\mathcal{C}))$. If there existed some nonzero $u\in\mathsf{K}(x)$, then by definition of $\mathsf{K}(x)$,
$u^{\mathsf{T}}(c-x)\leq 0,\qquad\forall c\in\mathcal{C}.$ (64)
By convexity, this would imply
$u^{\mathsf{T}}(y-x)\leq 0,\qquad\forall y\in\operatorname{conv}(\mathcal{C}).$ (65)
However, since $x$ is an interior point of $\operatorname{conv}(\mathcal{C})$, there exists $\varepsilon>0$ such that
$x+\varepsilon\frac{u}{\|u\|}\in\operatorname{conv}(\mathcal{C}),$ (66)
which would give
$u^{\mathsf{T}}\!\left(x+\varepsilon\frac{u}{\|u\|}-x\right)=\varepsilon\|u\|>0,$ (67)
a contradiction. Therefore $\mathsf{K}(x)$ contains no nonzero direction, so
$A_{x}=\mathsf{K}(x)\cap\mathbb{S}^{d-1}=\varnothing.$ (68)
This completes the proof. ∎
V Conclusion
This paper developed a geometric asymptotic perspective on one-shot detection under additive white Cauchy noise. Although isotropic Cauchy noise induces the same Euclidean Voronoi decision regions as Gaussian noise, the corresponding reliability laws differ fundamentally because the probability mass assigned to those regions is redistributed by the heavy-tailed noise geometry. In the small-noise regime, we established a reciprocal distance-spectrum upper bound showing that error probabilities retain sensitivity to the full constellation geometry. In the large-noise regime, we proved that the correct-decision probability converges to an angular limit determined solely by the Voronoi recession cone. Together, these results formalize a regime-dependent transition from distance-based to angle-based reliability descriptors. The worked examples in Appendices A and B show that this transition can lead either to highly nonuniform symbol reliability, including complete geometric collapse, or to symmetric limiting behavior in standard constellations such as 4QAM.
Appendix A Geometric Collapse: A Four-Point Example
To concretely illustrate Theorems 1 and 2, we analyze the four-point constellation
$\mathcal{X}=\{P_{1},P_{2},P_{3},P_{4}\}\subset\mathbb{R}^{2},$ (69)
with
$P_{1}$ $=(0,0),$ $P_{2}$ $=(1,0),$ (70)
$P_{3}$ $=(0,1),$ $P_{4}$ $=(0,-1).$
This example is deliberately minimal: all points share the same nearest-neighbor distance
$d_{\min}=1,$ (71)
yet their large-noise behavior under AWCN differs substantially because their Voronoi recession cones have different angular measures.
[…]
Figure 1: Voronoi geometry for the constellation $\mathcal{X}$ (see Appendix A). Left: Euclidean Voronoi cells. Right: recession cones and their angular patches on $\mathbb{S}^{1}$. The point $P_{1}$ has a degenerate cone of zero angular measure, whereas $P_{2},P_{3},P_{4}$ retain nontrivial patches. The nondegenerate patches correspond to $P_{2}$, $P_{3}$, and $P_{4}$, with opening angles $\pi/2$, $3\pi/4$, and $3\pi/4$, respectively.
The AWGN column is included only as a qualitative comparison: under Gaussian noise, the nearest-neighbor structure largely governs the high-reliability regime, whereas under AWCN the small-noise upper bound depends on the full reciprocal distance spectrum and the large-noise limit depends only on angular measures.
A-A Small-Noise Limit: Explicit Distance-Spectrum Evaluation
We first compute the pairwise distances:
$d_{12}=d_{13}=d_{14}$ $=1,$
$d_{23}=d_{24}$ $=\sqrt{2},$
$d_{34}$ $=2.$ (72)
By Theorem 1, for each transmitted symbol $P_{i}$,
$P_{e}(P_{i};\gamma)\leq\frac{2\gamma}{\pi}\sum_{j\neq i}\frac{1}{d_{ij}}+O(\gamma^{3}),\qquad\gamma\to 0.$ (73)
For $P_{1}=(0,0)$, all three competitors are at unit distance, so
$P_{e}(P_{1};\gamma)\leq\frac{2\gamma}{\pi}\left(1+1+1\right)+O(\gamma^{3})=\frac{6}{\pi}\gamma+O(\gamma^{3}).$ (74)
For $P_{2}=(1,0)$, the distances to $P_{1},P_{3},P_{4}$ are $1,\sqrt{2},\sqrt{2}$, respectively. Hence
$P_{e}(P_{2};\gamma)$ $\leq\frac{2\gamma}{\pi}\left(1+\frac{1}{\sqrt{2}}+\frac{1}{\sqrt{2}}\right)+O(\gamma^{3})$ (75)
$=\frac{2}{\pi}(1+\sqrt{2})\gamma+O(\gamma^{3}).$
For $P_{3}=(0,1)$, the distances to $P_{1},P_{2},P_{4}$ are $1,\sqrt{2},2$, respectively. Therefore
$P_{e}(P_{3};\gamma)$ $\leq\frac{2\gamma}{\pi}\left(1+\frac{1}{\sqrt{2}}+\frac{1}{2}\right)+O(\gamma^{3})$ (76)
$=\frac{1}{\pi}(3+\sqrt{2})\gamma+O(\gamma^{3}).$
By symmetry, the same expression holds for $P_{4}=(0,-1)$:
$P_{e}(P_{4};\gamma)\leq\frac{1}{\pi}(3+\sqrt{2})\gamma+O(\gamma^{3}).$ (77)
These calculations reproduce exactly the AWCN small-$\gamma$ column in Table I. In particular, although all four points share the same minimum distance, their reciprocal distance sums differ, so Theorem 1 already distinguishes them at leading order.
TABLE I: Quantitative transition from distance-based to angular descriptors for $\mathcal{X}$ (an illustrative example in Appendix A).
Constellation Point AWGN AWCN AWCN
$x$ (small $\sigma$) (small $\gamma$) ($\gamma\to\infty$)
$P_{1}=(0,0)$ $\approx 3\,Q\!\left(\frac{1}{2\sigma}\right)$ $\leq\frac{6}{\pi}\gamma+O(\gamma^{3})$ 11
$P_{2}=(1,0)$ $\approx Q\!\left(\frac{1}{2\sigma}\right)$ $\leq\frac{2}{\pi}(1+\sqrt{2})\gamma+O(\gamma^{3})$ $3/4$
$P_{3}=(0,1)$ $\approx Q\!\left(\frac{1}{2\sigma}\right)$ $\leq\frac{1}{\pi}(3+\sqrt{2})\gamma+O(\gamma^{3})$ $5/8$
$P_{4}=(0,-1)$ $\approx Q\!\left(\frac{1}{2\sigma}\right)$ $\leq\frac{1}{\pi}(3+\sqrt{2})\gamma+O(\gamma^{3})$ $5/8$
A-B Large-Noise Limit: Explicit Angular-Patch Evaluation
We now verify Theorem 2 by computing the recession cone
$\mathsf{K}(x)=\bigcap_{c\in\mathcal{X}\setminus\{x\}}\{u\in\mathbb{R}^{2}:\ u^{\mathsf{T}}(c-x)\leq 0\}$ (78)
for each constellation point. In two dimensions, Theorem 2 reduces to
$\lim_{\gamma\to\infty}P_{c}(x;\gamma)=\frac{\text{arc length of }A_{x}}{2\pi},\qquad A_{x}=\mathsf{K}(x)\cap\mathbb{S}^{1}.$ (79)
1) Point $P_{1}=(0,0)$
The competitor offsets are
$P_{2}-P_{1}$ $=(1,0),$ (80)
$P_{3}-P_{1}$ $=(0,1),$
$P_{4}-P_{1}$ $=(0,-1).$
Hence the cone constraints are
$u_{x}\leq 0,\quad u_{y}\leq 0,\quad-u_{y}\leq 0.$ (81)
Equivalently,
$u_{x}\leq 0,\quad u_{y}=0.$ (82)
Thus $\mathsf{K}(P_{1})$ is the negative $x$-axis ray, and the angular patch $A_{P_{1}}$ consists of a single point on $\mathbb{S}^{1}$. Its arc length is therefore zero, so
$\lim_{\gamma\to\infty}P_{c}(P_{1};\gamma)=0,\qquad\lim_{\gamma\to\infty}P_{e}(P_{1};\gamma)=1.$ (83)
This is the geometric-collapse phenomenon predicted by Corollary 1.
2) Point $P_{2}=(1,0)$
The competitor offsets are
$P_{1}-P_{2}$ $=(-1,0),$ (84)
$P_{3}-P_{2}$ $=(-1,1),$
$P_{4}-P_{2}$ $=(-1,-1).$
Hence
$u_{x}\leq 0,\quad-u_{x}+u_{y}\leq 0,\quad-u_{x}-u_{y}\leq 0,$ (85)
which is equivalent to
$u_{x}\geq 0,\quad|u_{y}|\leq u_{x}.$ (86)
Therefore $\mathsf{K}(P_{2})$ is a wedge centered on the positive $x$-axis with opening angle $\pi/2$, so its spherical patch has arc length $\pi/2$. By (79),
$\lim_{\gamma\to\infty}P_{c}(P_{2};\gamma)=\frac{\pi/2}{2\pi}=\frac{1}{4},\qquad\lim_{\gamma\to\infty}P_{e}(P_{2};\gamma)=\frac{3}{4}.$ (87)
3) Point $P_{3}=(0,1)$
The competitor offsets are
$P_{1}-P_{3}$ $=(0,-1),$ (88)
$P_{2}-P_{3}$ $=(1,-1),$
$P_{4}-P_{3}$ $=(0,-2).$
Hence
$u_{y}\leq 0,\quad u_{x}-u_{y}\leq 0,\quad-2u_{y}\leq 0,$ (89)
which reduces to
$u_{y}\geq 0,\quad u_{x}\leq u_{y}.$ (90)
Thus $\mathsf{K}(P_{3})$ is the wedge spanning angles from $\pi/4$ to $\pi$, whose opening angle is $3\pi/4$. Therefore
$\lim_{\gamma\to\infty}P_{c}(P_{3};\gamma)=\frac{3\pi/4}{2\pi}=\frac{3}{8},\qquad\lim_{\gamma\to\infty}P_{e}(P_{3};\gamma)=\frac{5}{8}.$ (91)
4) Point $P_{4}=(0,-1)$
By symmetry,
$P_{1}-P_{4}$ $=(0,1),$ (92)
$P_{2}-P_{4}$ $=(1,1),$
$P_{3}-P_{4}$ $=(0,2),$
so the cone constraints are
$u_{y}\leq 0,\quad u_{x}+u_{y}\leq 0.$ (93)
Equivalently,
$u_{y}\leq 0,\quad u_{x}\leq-u_{y}.$ (94)
Hence $\mathsf{K}(P_{4})$ is the reflected wedge spanning angles from $\pi$ to $7\pi/4$, again with opening angle $3\pi/4$. Thus
$\lim_{\gamma\to\infty}P_{c}(P_{4};\gamma)=\frac{3\pi/4}{2\pi}=\frac{3}{8},\qquad\lim_{\gamma\to\infty}P_{e}(P_{4};\gamma)=\frac{5}{8}.$ (95)
Under equiprobable signaling, symmetry implies that the average correct-decision probability also satisfies
$\lim_{\gamma\to\infty}P_{c}(\gamma)$ $=\frac{1}{4}\left(0+\frac{1}{4}+\frac{3}{8}+\frac{3}{8}\right)$ (96)
$=\frac{1}{4},$
and hence
$\lim_{\gamma\to\infty}P_{e}(\gamma)=1-\frac{1}{4}=\frac{3}{4}.$ (97)
This example shows explicitly that the small-noise and large-noise regimes are governed by fundamentally different geometric descriptors. In the high-reliability regime, the relevant quantity is the reciprocal distance sum appearing in Theorem 1; in the infinite-noise regime, all bounded distance information disappears and only the angular patch measure from Theorem 2 survives.
Appendix B A Standard 4QAM Sanity Check
As a complementary sanity check, we briefly examine the standard 44QAM constellation
$\mathcal{Q}=\{Q_{1},Q_{2},Q_{3},Q_{4}\}\subset\mathbb{R}^{2},$ (98)
with
$Q_{1}$ $=(1,1),$ $Q_{2}$ $=(-1,1),$ (99)
$Q_{3}$ $=(-1,-1),$ $Q_{4}$ $=(1,-1).$
Unlike the asymmetric four-point example in Appendix A, this constellation is fully symmetric. As a result, both the small-noise distance descriptor and the large-noise angular descriptor are identical for all four symbols.
B-A Small-Noise Regime
For any fixed point, say $Q_{1}=(1,1)$, the distances to the other three points are
$\|Q_{1}-Q_{2}\|=2,\quad\|Q_{1}-Q_{4}\|=2,\quad\|Q_{1}-Q_{3}\|=2\sqrt{2}.$ (100)
Hence, by Theorem 1,
$P_{e}(Q_{1};\gamma)$ $\leq\frac{2\gamma}{\pi}\left(\frac{1}{2}+\frac{1}{2}+\frac{1}{2\sqrt{2}}\right)+O(\gamma^{3})$ (101)
$=\frac{1}{\pi}\left(2+\frac{1}{\sqrt{2}}\right)\gamma+O(\gamma^{3}),\quad\gamma\to 0.$
By symmetry, the same bound holds for $Q_{2},Q_{3},Q_{4}$. Thus all symbols have the same leading-order small-noise behavior.
B-B Large-Noise Regime
We next evaluate the recession cones. For $Q_{1}=(1,1)$, the competitor offsets are
$Q_{2}-Q_{1}$ $=(-2,0),$ (102)
$Q_{3}-Q_{1}$ $=(-2,-2),$
$Q_{4}-Q_{1}$ $=(0,-2).$
Therefore, the cone constraints are
$2u_{x}\leq 0,\quad-2u_{x}-2u_{y}\leq 0,\quad-2u_{y}\leq 0,$ (103)
which implies
$u_{x}\geq 0,\quad u_{y}\geq 0.$ (104)
Hence $\mathsf{K}(Q_{1})$ is the first quadrant, whose angular patch on $\mathbb{S}^{1}$ has arc length $\pi/2$. By Theorem 2,
$\lim_{\gamma\to\infty}P_{c}(Q_{1};\gamma)=\frac{\pi/2}{2\pi}=\frac{1}{4},\quad\lim_{\gamma\to\infty}P_{e}(Q_{1};\gamma)=\frac{3}{4}.$ (105)
Again by symmetry, the same limit holds for all four symbols:
$\lim_{\gamma\to\infty}P_{c}(Q_{i};\gamma)=\frac{1}{4},\quad\lim_{\gamma\to\infty}P_{e}(Q_{i};\gamma)=\frac{3}{4},$ (106)
for $i=1,2,3,4$. Under equiprobable signaling, the average correct-decision probability therefore satisfies
$\lim_{\gamma\to\infty}P_{c}(\gamma)=\frac{1}{4},\quad\lim_{\gamma\to\infty}P_{e}(\gamma)=\frac{3}{4}.$ (107)
This example serves as a useful baseline: for a highly symmetric constellation such as 44QAM, both the reciprocal distance sum in Theorem 1 and the angular patch measure in Theorem 2 are identical across symbols. In contrast, the asymmetric example in Appendix A shows that these descriptors can vary substantially from point to point, leading to nonuniform reliability in both the small-noise and large-noise regimes.