### 20 Apr 2026 v2

# From Distance to Angle: One-Shot Detection Under Isotropic Multivariate Cauchy Noise

Abstract
We study one-shot detection under isotropic multivariate Cauchy noise using finite constellations, with emphasis on the geometric mechanisms governing symbol-level reliability. Under isotropic Cauchy noise, the maximum-likelihood rule induces the same Euclidean Voronoi decision regions as in the Gaussian case, so the distinction lies not in the decision geometry itself but in how probability mass is distributed over these fixed regions. In the small-noise regime, we derive a reciprocal distance-spectrum upper bound for the symbol error probability, showing that reliability retains a longer-range dependence on the global constellation geometry than under additive white Gaussian noise. In the large-noise regime, we prove that the correct-decision probability converges to a limit determined solely by the angular measure of the associated Voronoi recession cone. These results formalize a regime-dependent transition from distance-based to angle-based reliability descriptors under heavy-tailed noise. The theory is further illustrated through an asymmetric four-point example exhibiting geometric collapse, a standard 4QAM sanity check, and finite-$\gamma$ numerical validation for both asymptotic regimes.
I Introduction
Additive white Gaussian noise (AWGN) has long served as the canonical model for detection and constellation design, under which reliability is largely governed by minimum-distance geometry. This viewpoint underlies much of modern digital communication theory [13], particularly in the small-noise regime. Recent studies, however, have identified communication and sensing scenarios in which noise exhibits heavy-tailed, non-Gaussian behavior, rendering Gaussian-based design principles suboptimal [14, 10]. Such heavy-tailed noise models have also attracted sustained interest across information theory, signal processing, and communication systems [2, 16, 12].
Specifically, the Cauchy model has emerged as a prototypical member of the $\alpha$-stable family [10], offering a mathematically rigorous yet tractable departure from the Gaussian assumption. While its infinite variance and polynomial tail decay complicate traditional concentration-based analysis (e.g., via the Law-of-Large-Numbers (LLN)) [16], the model is far from a mere theoretical abstraction. It arises naturally in practical systems, including massive MIMO [7] and diffusion-based molecular communication [5, 8]. At a broader level, existing work on stable and Cauchy noise has progressed along several largely separate directions. In information theory, the adopted input constraint often changes with the characteristic exponent: the Gaussian case is classically studied under a second-moment constraint, the Cauchy case has been analyzed under logarithmic cost [3], and more general $\alpha$-stable channels have also been treated under alternative moment constraints to obtain tractable capacity bounds [1]. In signal processing, impulsive environments have motivated generalized information and estimation tools beyond the Gaussian second-order framework [4]. In communication systems, heavy-tailed laws also arise from concrete channel models rather than from abstract distributional choices alone; in particular, Cauchy-type behavior can naturally emerge in molecular communication settings such as first-arrival-position channels [8, 9].
These developments point to a common missing piece. Although recent findings [12] show that memoryless scalar Cauchy-noise channels can sustain positive information rates under a second-moment input (energy) constraint, and thus support a renewed power-constrained viewpoint in the Cauchy regime, a multivariate geometric theory for finite-constellation detection under isotropic multivariate Cauchy noise remains undeveloped. In particular, from the perspective of symbol-level detection, it is still unclear which geometric descriptors replace the classical minimum-distance paradigm once the noise law becomes heavy-tailed. This motivates a timely re-examination of finite-constellation detection and design (cf. [15, 13]) in the presence of Cauchy noise.
To address this need, this paper develops a geometric asymptotic theory for one-shot detection over channels corrupted by isotropic multivariate Cauchy noise. Our main contributions are summarized as follows:
  • We formulate one-shot finite-constellation detection under isotropic multivariate Cauchy noise and clarify that, although the maximum-likelihood rule induces the same Euclidean Voronoi decision regions as in the Gaussian case, the associated reliability laws are governed by fundamentally different geometric mechanisms.
  • In the small-noise regime, we derive a reciprocal distance-spectrum upper bound showing that one-shot reliability retains a longer-range sensitivity to the global constellation geometry than in the Gaussian setting.
  • In the large-noise regime, we prove that bounded distance information disappears asymptotically and that the correct-decision probability converges to a limit determined solely by the angular measure of the associated Voronoi recession cone (cf. [11]), thereby identifying an angle-based reliability descriptor and a corresponding geometric-collapse criterion.
  • We further support this regime-dependent transition through explicit worked examples and finite-$\gamma$ numerical validation, showing how the theory manifests in both asymmetric (see Appendix A) and symmetric (see Appendix B) constellations.
The remainder of this paper is organized as follows. Section II introduces the system model and problem statement. Section III presents the small-noise distance-spectrum bound. Section IV establishes the large-noise angular convergence theorem and the associated geometric-collapse criterion. Section V provides finite-$\gamma$ numerical validation for both asymptotic regimes. The main results are further illustrated through an asymmetric four-point worked example in Appendix A and a standard 44QAM sanity check in Appendix B. Conclusions are drawn in Section VI.
II System Model and Problem Statement
II-A Vector Cauchy Channels
We focus on one-shot detection (also referred to as symbol-by-symbol detection), where the receiver performs optimal maximum-likelihood (ML) decoding based solely on the current observation vector. This setting isolates the symbol-level effect of heavy-tailed noise on decision geometry and reliability, without invoking coding across multiple channel uses.
We consider a $d$-dimensional discrete-time vector channel model defined as
$\bm{y}=\bm{x}+\bm{n},$ (1)
where $\bm{x}\in\mathcal{C}\subset\mathbb{R}^{d}$ is the transmitted signal vector from a finite constellation subject to an average energy constraint $\mathbb{E}[\|\bm{x}\|^{2}]\leq P$. The additive noise $\bm{n}\in\mathbb{R}^{d}$ is modeled as an isotropic $d$-dimensional Cauchy random vector with probability density function (PDF) (cf. [14, 16]),
$f_{\bm{N}}(\bm{n})=c_{d}\,\gamma^{-d}\left(1+\frac{\|\bm{n}\|^{2}}{\gamma^{2}}\right)^{-\frac{d+1}{2}},\quad c_{d}=\frac{\Gamma\!\left(\frac{d+1}{2}\right)}{\pi^{\frac{d+1}{2}}},$ (2)
where $\gamma>0$ is the scale parameter controlling the distribution’s width. Here, the noise is modeled as a single isotropic multivariate Cauchy vector in $\mathbb{R}^{d}$, rather than as independent componentwise Cauchy coordinates. Unlike Gaussian noise, the Cauchy distribution lacks finite moments; thus, standard signal-to-noise ratio (SNR) is technically undefined. We instead use the normalized proxy $P/\gamma^{2}$ to characterize system performance (cf. [12]).
In the Gaussian setting, isotropy and componentwise i.i.d. structure lead to the same multivariate law. In the Cauchy setting, however, these two notions are no longer equivalent: the isotropic multivariate Cauchy law is radially symmetric, whereas independent Cauchy coordinates define a different joint distribution. This distinction matters here because our analysis is carried out under an isotropic multivariate Cauchy observation model rather than a componentwise independent one. It also places the present work in contrast with the memoryless scalar Cauchy-noise setting studied in [12].
Under this spherically symmetric model, the ML decoding rule minimizes the negative log-likelihood, which reduces to the standard Euclidean distance metric. Namely,
$\hat{\bm{x}}=\operatorname*{arg\,min}_{\bm{c}\in\mathcal{C}}\log\left(1+\frac{\|\bm{y}-\bm{c}\|^{2}}{\gamma^{2}}\right)\equiv\operatorname*{arg\,min}_{\bm{c}\in\mathcal{C}}\|\bm{y}-\bm{c}\|.$ (3)
Because the isotropic Cauchy density is monotonically decreasing with the Euclidean norm, the exact decision boundaries induced by (3) correspond to Euclidean Voronoi cells (cf. [11]), entirely independent of $\gamma$. Under equiprobable signaling, the average symbol error probability is formally given by
$P_{e}=1-\frac{1}{M}\sum_{\bm{x}\in\mathcal{C}}\int_{V(\bm{x})}f_{\bm{N}}(\bm{y}-\bm{x})~\mathrm{d}\bm{y}.$ (4)
For the rest of this paper, we use uppercase letters such as $X,Y,N$ for random variables and lowercase letters such as $x,y,n$ for their realizations.
II-B The Combining Gain Paradox
In traditional AWGN vector channels, linear aggregation strategies (e.g., maximum ratio combining) often yield an averaging gain under standard finite-variance assumptions, so reliability can improve with the number of spatial dimensions or channel paths [13]. Under isotropic multivariate Cauchy noise, this self-averaging intuition breaks down: the Cauchy law has neither finite mean nor finite variance, so the usual LLN viewpoint no longer applies. More fundamentally, under the isotropic multivariate Cauchy model, every one-dimensional projection onto a unit direction remains a scalar Cauchy random variable with the same scale parameter (cf. [14, 10]). Consequently, linear combining does not reduce the effective noise scale in the usual Gaussian sense, but instead produces another scalar Cauchy projection, thereby removing the classical concentration effect associated with multidimensional combining.
Because the exact $d$-dimensional Cauchy integral in (4) lacks a closed-form solution over arbitrary Voronoi polytopes, understanding this contrast calls for a viewpoint beyond linear averaging arguments alone. The specific objective of this paper is therefore to geometrically reinterpret the error integral (4). By showing how symbol reliability transitions from distance-based to angle-based behavior, we aim to identify the geometric descriptors that govern one-shot detection under the isotropic Cauchy model and that may inform future geometry-aware receiver design.
III Small-Noise Regime: Distance Spectrum Bounds
In this section, we formally characterize the one-shot reliability in the small-noise regime.
III-A Pairwise Comparison Term
Before deriving the small-noise bound, we isolate the basic pairwise comparison event that enters the union-bound argument. Fix two distinct constellation points $x_{i},x_{j}\in\mathcal{C}$, and define
$d_{ij}:=\|x_{i}-x_{j}\|,\qquad u_{ij}:=\frac{x_{j}-x_{i}}{\|x_{j}-x_{i}\|}.$ (5)
Assume that $x_{i}$ is transmitted. Since the isotropic Cauchy law is absolutely continuous, tie events occur with probability zero and do not affect the ML analysis. Under ML decoding, the event that $x_{j}$ is at least as favorable as $x_{i}$ is
$\|Y-x_{j}\|\leq\|Y-x_{i}\|.$ (6)
Since $Y=x_{i}+N$, this is equivalent to
$\|N-(x_{j}-x_{i})\|\leq\|N\|,$ (7)
or, after expanding the squared norms,
$u_{ij}^{\mathsf{T}}N\geq\frac{d_{ij}}{2}.$ (8)
Because the noise vector $N$ is isotropic Cauchy, every one-dimensional projection onto a unit direction remains a scalar Cauchy random variable with the same scale parameter $\gamma$. Therefore,
$\mathbb{P}\!\left(\|Y-x_{j}\|\leq\|Y-x_{i}\|\,\middle|\,X=x_{i}\right)$ $=\mathbb{P}\!\left(u_{ij}^{\mathsf{T}}N\geq\frac{d_{ij}}{2}\right)$ (9)
$=\frac{1}{2}-\frac{1}{\pi}\arctan\!\left(\frac{d_{ij}}{2\gamma}\right).$
Equation (9) shows that each pairwise comparison term depends only on the Euclidean distance $d_{ij}$. This quantity will serve as the basic building block in the distance-spectrum bound of Theorem 1.
III-B Main Result: Small-Noise Bound and Asymptotics
We define the distance spectrum of $\mathcal{C}$ as the multiset $\mathcal{S}(\mathcal{C})=\{d_{ij}:1\leq i<j\leq M\}$. The following theorem establishes the asymptotic behavior of the error probability as the noise scale vanishes.
Theorem 1 (Small-Noise Distance-Spectrum Bound).
Let $\mathcal{C}=\{x_{1},\dots,x_{M}\}\subset\mathbb{R}^{d}$ be a finite constellation, and consider the isotropic $d$-dimensional Cauchy channel
$Y=X+N,$ (10)
with maximum-likelihood decoding as in (3). For a fixed transmitted symbol $x_{i}\in\mathcal{C}$, define
$d_{ij}:=\|x_{i}-x_{j}\|,\qquad j\neq i.$ (11)
Then, for every $\gamma>0$, the conditional symbol error probability satisfies
$P_{e}(x_{i};\gamma)\leq\sum_{j\neq i}\left[\frac{1}{2}-\frac{1}{\pi}\arctan\!\left(\frac{d_{ij}}{2\gamma}\right)\right].$ (12)
Consequently, as $\gamma\to 0$,
$P_{e}(x_{i};\gamma)\leq\frac{2\gamma}{\pi}\sum_{j\neq i}\frac{1}{d_{ij}}+O(\gamma^{3}).$ (13)
If the symbols are equiprobable, then the average symbol error probability satisfies
$P_{e}(\gamma)\leq\frac{4\gamma}{M\pi}\sum_{1\leq i<j\leq M}\frac{1}{d_{ij}}+O(\gamma^{3}).$ (14)
Hence, in the small-noise regime, the one-shot reliability is controlled, at leading order, by the reciprocal distance spectrum of the constellation.
Proof.
Fix $x_{i}\in\mathcal{C}$ and assume that $x_{i}$ is transmitted. Let
$E_{i}:=\{\hat{X}\neq x_{i}\}$ (15)
denote the decoding error event under ML decoding.
For each competitor $x_{j}\in\mathcal{C}\setminus\{x_{i}\}$, define
$H_{ij}:=\Bigl\{n\in\mathbb{R}^{d}:\|x_{i}+n-x_{j}\|\leq\|n\|\Bigr\}.$ (16)
The event $H_{ij}$ means that, after adding noise $n$, the point $x_{j}$ is at least as close to the received vector as $x_{i}$. Therefore,
$E_{i}\subseteq\bigcup_{j\neq i}H_{ij},$ (17)
and hence, by the union bound,
$P_{e}(x_{i};\gamma)\leq\sum_{j\neq i}\mathbb{P}(H_{ij}).$ (18)
We now simplify each event $H_{ij}$. Expanding the squared norms gives
$\|x_{i}+n-x_{j}\|^{2}\leq\|n\|^{2}$ (19)
$\Longleftrightarrow\quad\|n-(x_{j}-x_{i})\|^{2}\leq\|n\|^{2},$
which is equivalent to
$2(x_{j}-x_{i})^{\mathsf{T}}n+\|x_{j}-x_{i}\|^{2}\leq 0.$ (20)
Thus,
$H_{ij}=\left\{(x_{j}-x_{i})^{\mathsf{T}}N\geq\frac{\|x_{j}-x_{i}\|^{2}}{2}\right\}.$ (21)
Let
$u_{ij}:=\frac{x_{j}-x_{i}}{\|x_{j}-x_{i}\|}=\frac{x_{j}-x_{i}}{d_{ij}}.$ (22)
Then
$H_{ij}=\left\{u_{ij}^{\mathsf{T}}N\geq\frac{d_{ij}}{2}\right\}.$ (23)
Because $N$ is an isotropic multivariate Cauchy vector with scale parameter $\gamma$, every one-dimensional projection $u^{\mathsf{T}}N$ onto a unit vector $u$ is a one-dimensional Cauchy random variable with location 0 and scale $\gamma$, i.e.,
$u^{\mathsf{T}}N\sim\mathrm{Cauchy}(0,\gamma),$ (24)
whose density is
$f_{Z}(z)=\frac{1}{\pi\gamma}\frac{1}{1+(z/\gamma)^{2}},\qquad z\in\mathbb{R}.$ (25)
Therefore,
$\mathbb{P}(H_{ij})$ $=\mathbb{P}\!\left(u_{ij}^{\mathsf{T}}N\geq\frac{d_{ij}}{2}\right)$ (26)
$=\int_{d_{ij}/2}^{\infty}\frac{1}{\pi\gamma}\frac{1}{1+(z/\gamma)^{2}}~\mathrm{d}z.$
Evaluating the integral yields
$\mathbb{P}(H_{ij})=\frac{1}{2}-\frac{1}{\pi}\arctan\!\left(\frac{d_{ij}}{2\gamma}\right).$ (27)
Substituting this into (18) proves (12).
To obtain the small-noise expansion, use the standard asymptotic formula [6]:
$\arctan t=\frac{\pi}{2}-\frac{1}{t}+\frac{1}{3t^{3}}+O(t^{-5}),\qquad t\to\infty.$ (28)
Setting $t=d_{ij}/(2\gamma)$ gives, for each fixed $i\neq j$,
$\frac{1}{2}-\frac{1}{\pi}\arctan\!\left(\frac{d_{ij}}{2\gamma}\right)=\frac{2\gamma}{\pi d_{ij}}+O(\gamma^{3}),\qquad\gamma\to 0.$ (29)
Since the constellation is finite, summing over $j\neq i$ yields
$P_{e}(x_{i};\gamma)\leq\frac{2\gamma}{\pi}\sum_{j\neq i}\frac{1}{d_{ij}}+O(\gamma^{3}),$ (30)
which proves (13).
Finally, if the input symbols are equiprobable, then
$P_{e}(\gamma)=\frac{1}{M}\sum_{i=1}^{M}P_{e}(x_{i};\gamma).$ (31)
Averaging the previous bound over $i$ gives
$P_{e}(\gamma)\leq\frac{2\gamma}{M\pi}\sum_{i=1}^{M}\sum_{j\neq i}\frac{1}{d_{ij}}+O(\gamma^{3}).$ (32)
Using the symmetry $d_{ij}=d_{ji}$ and double counting,
$\sum_{i=1}^{M}\sum_{j\neq i}\frac{1}{d_{ij}}=2\sum_{1\leq i<j\leq M}\frac{1}{d_{ij}},$ (33)
which proves (14). ∎
Remark.
Theorem 1 clarifies an important contrast between Gaussian and Cauchy detection. Under AWGN, pairwise error terms decay exponentially with distance, so the nearest-neighbor structure typically dominates the small-noise behavior. Under the isotropic Cauchy model, however, the pairwise comparison term in (9) satisfies
$\frac{1}{2}-\frac{1}{\pi}\arctan\!\left(\frac{d_{ij}}{2\gamma}\right)=\frac{2\gamma}{\pi d_{ij}}+O(\gamma^{3}),\qquad\gamma\to 0,$ (34)
which decays only algebraically in the distance $d_{ij}$. As a result, farther constellation points are suppressed more weakly than in the Gaussian case, and the small-noise error bound naturally depends on the full reciprocal distance spectrum rather than solely on the minimum distance. In this sense, Theorem 1 should be interpreted as a geometric upper bound showing that, even in the high-reliability regime, Cauchy noise preserves a longer-range sensitivity to the global constellation geometry.
IV Large-Noise Regime: Angular Voronoi Convergence
As $\gamma$ increases, the distance-based description breaks down. In this regime, the received vector is pushed far away from the transmitted symbol, so the bounded offsets among constellation points become negligible relative to the overall noise magnitude. We show that the decoding problem then loses sensitivity to finite Euclidean distances and retains only the asymptotic angular geometry of the Voronoi cells. The appropriate descriptor is therefore the recession cone of each Voronoi region and its associated spherical angular patch.
IV-A Voronoi Cones and Angular Patches
For each $x\in\mathcal{C}$, the associated Voronoi cell is
$V(x)=\{y\in\mathbb{R}^{d}:\ \|y-x\|\leq\|y-c\|,\ \forall c\in\mathcal{C}\}.$ (35)
Its recession cone is defined by
$\mathsf{K}(x)=\bigcap_{c\in\mathcal{C}\setminus\{x\}}\{u\in\mathbb{R}^{d}:\ u^{\mathsf{T}}(c-x)\leq 0\}.$ (36)
The spherical cross-section
$A_{x}:=\mathsf{K}(x)\cap\mathbb{S}^{d-1}$ (37)
is called the Voronoi angular patch of $x$, where $\mathbb{S}^{d-1}$ denotes the unit sphere in $\mathbb{R}^{d}$.
The geometric meaning of (36) is transparent after translating the Voronoi cell to the origin. Indeed,
$V(x)-x=\bigcap_{c\in\mathcal{C}\setminus\{x\}}\left\{z\in\mathbb{R}^{d}:\ z^{\mathsf{T}}(c-x)\leq\frac{\|c-x\|^{2}}{2}\right\}.$ (38)
Hence, after scaling by $\gamma$,
$\frac{V(x)-x}{\gamma}=\bigcap_{c\in\mathcal{C}\setminus\{x\}}\left\{u\in\mathbb{R}^{d}:\ u^{\mathsf{T}}(c-x)\leq\frac{\|c-x\|^{2}}{2\gamma}\right\},$ (39)
which suggests the recession-cone limit (36) as $\gamma\to\infty$. Thus, the Voronoi cone $\mathsf{K}(x)$ is precisely the large-noise limit of the translated-and-rescaled decision region.
IV-B Main Result: Large-Noise Convergence
The following theorem shows that, in the infinite-noise limit, the correct-decision probability depends only on the angular measure of the recession cone.
Theorem 2 (Large-Noise Convergence).
Let $\mathcal{C}\subset\mathbb{R}^{d}$ be a finite constellation, and let
$P_{c}(x;\gamma):=\mathbb{P}(\hat{X}=x\,|\,X=x)$ (40)
denote the conditional correct-decision probability under the isotropic $d$-dimensional Cauchy channel with scale parameter $\gamma$. For each $x\in\mathcal{C}$, let $\mathsf{K}(x)$ and $A_{x}$ be defined by (36) and (37), and let $\sigma_{d-1}$ denote the usual surface measure on $\mathbb{S}^{d-1}$. Then
$\lim_{\gamma\to\infty}P_{c}(x;\gamma)=\frac{\sigma_{d-1}(A_{x})}{\sigma_{d-1}(\mathbb{S}^{d-1})}.$ (41)
Equivalently,
$\lim_{\gamma\to\infty}P_{e}(x;\gamma)=1-\frac{\sigma_{d-1}(A_{x})}{\sigma_{d-1}(\mathbb{S}^{d-1})},$ (42)
where $P_{e}(x;\gamma)=1-P_{c}(x;\gamma)$.
If the constellation symbols are equiprobable, define
$P_{c}(\gamma):=\frac{1}{M}\sum_{x\in\mathcal{C}}P_{c}(x;\gamma).$ (43)
Then
$\lim_{\gamma\to\infty}P_{c}(\gamma)=\frac{1}{M}\sum_{x\in\mathcal{C}}\frac{\sigma_{d-1}(A_{x})}{\sigma_{d-1}(\mathbb{S}^{d-1})}.$ (44)
Proof.
Fix $x\in\mathcal{C}$. By definition of ML decoding,
$P_{c}(x;\gamma)=\mathbb{P}(Y\in V(x)\mid X=x)=\mathbb{P}(N\in V(x)-x).$ (45)
Since the noise density is
$f_{N}(n)=c_{d}\,\gamma^{-d}\left(1+\frac{\|n\|^{2}}{\gamma^{2}}\right)^{-\frac{d+1}{2}},$ (46)
we obtain
$P_{c}(x;\gamma)=\int_{V(x)-x}c_{d}\,\gamma^{-d}\left(1+\frac{\|n\|^{2}}{\gamma^{2}}\right)^{-\frac{d+1}{2}}~\mathrm{d}n.$ (47)
Applying the change of variables $n=\gamma u$ yields
$P_{c}(x;\gamma)=\int_{(V(x)-x)/\gamma}c_{d}\,(1+\|u\|^{2})^{-\frac{d+1}{2}}~\mathrm{d}u.$ (48)
We now show that the indicator of the scaled region converges almost everywhere to the indicator of the recession cone. From (39),
$\frac{V(x)-x}{\gamma}=\bigcap_{c\in\mathcal{C}\setminus\{x\}}\left\{u:\ u^{\mathsf{T}}(c-x)\leq\frac{\|c-x\|^{2}}{2\gamma}\right\}.$ (49)
Let
$B_{x}:=\bigcup_{c\in\mathcal{C}\setminus\{x\}}\{u\in\mathbb{R}^{d}:\ u^{\mathsf{T}}(c-x)=0\}.$ (50)
Since $\mathcal{C}$ is finite, $B_{x}$ is a finite union of hyperplanes and therefore has Lebesgue measure zero.
Now fix $u\notin B_{x}$. Then for every $c\neq x$, the quantity $u^{\mathsf{T}}(c-x)$ is nonzero. If $u\in\mathsf{K}(x)$, then all these inner products are strictly negative, so for sufficiently large $\gamma$ they satisfy
$u^{\mathsf{T}}(c-x)\leq\frac{\|c-x\|^{2}}{2\gamma},\qquad\forall c\neq x,$ (51)
and hence $u\in(V(x)-x)/\gamma$. Conversely, if $u\notin\mathsf{K}(x)$, then there exists some $c\neq x$ such that $u^{\mathsf{T}}(c-x)>0$, and for sufficiently large $\gamma$,
$u^{\mathsf{T}}(c-x)>\frac{\|c-x\|^{2}}{2\gamma},$ (52)
so $u\notin(V(x)-x)/\gamma$. Therefore,
$\mathbf{1}_{(V(x)-x)/\gamma}(u)\to\mathbf{1}_{\mathsf{K}(x)}(u)\qquad\text{for a.e. }u\in\mathbb{R}^{d}.$ (53)
Since
$0\leq\mathbf{1}_{(V(x)-x)/\gamma}(u)\,c_{d}(1+\|u\|^{2})^{-\frac{d+1}{2}}\leq c_{d}(1+\|u\|^{2})^{-\frac{d+1}{2}},$ (54)
and the dominating function on the right is integrable over $\mathbb{R}^{d}$, the dominated convergence theorem applied to (48) and (53) gives
$\lim_{\gamma\to\infty}P_{c}(x;\gamma)=\int_{\mathsf{K}(x)}c_{d}(1+\|u\|^{2})^{-\frac{d+1}{2}}~\mathrm{d}u.$ (55)
It remains to evaluate the right-hand side in angular form. Because the density is radial, we use polar coordinates $u=r\theta$, where $r\in[0,\infty)$ and $\theta\in\mathbb{S}^{d-1}$. Since $\mathsf{K}(x)$ is a cone, its polar decomposition is exactly
$\mathsf{K}(x)=\{r\theta:\ r\geq 0,\ \theta\in A_{x}\}.$ (56)
Hence
$\int_{\mathsf{K}(x)}c_{d}(1+\|u\|^{2})^{-\frac{d+1}{2}}\mathrm{d}u$ (57)
$=\int_{A_{x}}\!\int_{0}^{\infty}c_{d}(1+r^{2})^{-\frac{d+1}{2}}r^{d-1}\mathrm{d}r\mathrm{d}\sigma_{d-1}(\theta)$
$=\sigma_{d-1}(A_{x})\underbrace{\int_{0}^{\infty}c_{d}(1+r^{2})^{-\frac{d+1}{2}}r^{d-1}\mathrm{d}r}_{=:I_{d}}.$
On the other hand, integrating the same radial density over the entire space $\mathbb{R}^{d}$ gives
$1=\int_{\mathbb{R}^{d}}c_{d}(1+\|u\|^{2})^{-\frac{d+1}{2}}~\mathrm{d}u=\sigma_{d-1}(\mathbb{S}^{d-1})\,I_{d}.$ (58)
Thus
$I_{d}=\frac{1}{\sigma_{d-1}(\mathbb{S}^{d-1})}.$ (59)
Substituting this into (57) and then into (55) yields
$\lim_{\gamma\to\infty}P_{c}(x;\gamma)=\frac{\sigma_{d-1}(A_{x})}{\sigma_{d-1}(\mathbb{S}^{d-1})},$ (60)
which proves (41). Equation (42) follows from $P_{e}(x;\gamma)=1-P_{c}(x;\gamma)$, and (44) follows by averaging over equiprobable symbols. ∎
(a) Average symbol error probability in the small-$\gamma$ regime for the asymmetric four-point constellation in Appendix A. The Monte Carlo estimate is compared with the exact upper bound from Theorem 1 and with its leading-order asymptotic approximation. Over the displayed range in (71), the Monte Carlo curve remains below the exact bound, while the exact bound and the leading-order asymptotic are nearly indistinguishable. This shows that the reciprocal-distance-spectrum scaling predicted by Theorem 1 becomes visible already at finite small noise scales.
(b) Conditional symbol error probabilities in the same small-$\gamma$ regime for the asymmetric four-point constellation in Appendix A. For each transmitted symbol, the Monte Carlo estimates for $P_{1},P_{2},P_{3},$ and $P_{4}$ are shown together with the corresponding exact conditional bound and leading-order asymptotic line from Theorem 1. The ordering predicted in Appendix A from the reciprocal-distance sums is clearly observed: $P_{1}$ is the least reliable symbol, followed by $P_{2}$, whereas $P_{3}$ and $P_{4}$ are more reliable and nearly coincide. Because all four points have the same minimum distance, this figure highlights that under isotropic Cauchy noise, symbol reliability depends on the full reciprocal-distance geometry rather than on $d_{\min}$ alone.
Figure 1: Finite-$\gamma$ validation of the small-noise theory in Theorem 1 for the asymmetric four-point constellation from Appendix A. Panel (a) verifies that the average Monte Carlo symbol error probability remains below the exact bound in Theorem 1 and that the leading-order asymptotic already tracks the bound closely in the finite small-noise regime. Panel (b) shows that the same phenomenon persists at the symbol level and that the conditional reliability ordering is governed by the reciprocal-distance coefficients derived in Theorem 1.
An immediate and striking consequence of Theorem 2 is the identification of constellation points that become entirely unreliable.
Corollary 1 (Condition for Geometric Collapse).
Let $x\in\mathcal{C}$. If the angular patch $A_{x}$ has zero surface measure, i.e.,
$\sigma_{d-1}(A_{x})=0,$ (61)
then
$\lim_{\gamma\to\infty}P_{c}(x;\gamma)=0,\qquad\lim_{\gamma\to\infty}P_{e}(x;\gamma)=1.$ (62)
In particular, if $x$ lies in the interior of $\operatorname{conv}(\mathcal{C})$, then $A_{x}=\varnothing$, and therefore geometric collapse occurs.
Proof.
If $\sigma_{d-1}(A_{x})=0$, then (41) in Theorem 2 immediately gives
$\lim_{\gamma\to\infty}P_{c}(x;\gamma)=\frac{\sigma_{d-1}(A_{x})}{\sigma_{d-1}(\mathbb{S}^{d-1})}=0,$ (63)
which proves (62).
For the interior-point claim, suppose that $x\in\operatorname{int}(\operatorname{conv}(\mathcal{C}))$. If there existed some nonzero $u\in\mathsf{K}(x)$, then by definition of $\mathsf{K}(x)$,
$u^{\mathsf{T}}(c-x)\leq 0,\qquad\forall c\in\mathcal{C}.$ (64)
By convexity, this would imply
$u^{\mathsf{T}}(y-x)\leq 0,\qquad\forall y\in\operatorname{conv}(\mathcal{C}).$ (65)
However, since $x$ is an interior point of $\operatorname{conv}(\mathcal{C})$, there exists $\varepsilon>0$ such that
$x+\varepsilon\frac{u}{\|u\|}\in\operatorname{conv}(\mathcal{C}),$ (66)
which would give
$u^{\mathsf{T}}\!\left(x+\varepsilon\frac{u}{\|u\|}-x\right)=\varepsilon\|u\|>0,$ (67)
a contradiction. Therefore $\mathsf{K}(x)$ contains no nonzero direction, so
$A_{x}=\mathsf{K}(x)\cap\mathbb{S}^{d-1}=\varnothing.$ (68)
This completes the proof. ∎
V Numerical Results
Unless otherwise stated, all Monte Carlo (MC) experiments in this section use equiprobable signaling, a fixed random seed rng(2026,’twister’), and $N_{\mathrm{mc}}=5\times 10^{5}$ trials for each value of $\gamma$, processed in batches of size $10^{5}$. The additive noise is generated as an isotropic multivariate Cauchy vector in $\mathbb{R}^{2}$ via
$N=\gamma\,\frac{G}{H},$ (69)
where $G\sim\mathcal{N}(0,I_{2})$ and $H\sim\mathcal{N}(0,1)$ are independent. In all simulations, maximum-likelihood decoding is implemented by the equivalent Euclidean nearest-neighbor rule in (3).
V-A Finite-$\gamma$ Validation of Theorem 1
We numerically validate Theorem 1 in the finite small-noise regime using the asymmetric four-point constellation from Appendix A,
$\mathcal{X}=\{(0,0),(1,0),(0,1),(0,-1)\}.$ (70)
For each
[…]
we estimate the symbol error probability by Monte Carlo simulation and compare it with both the exact union bound from Theorem 1 and its leading-order small-$\gamma$ asymptotic expression.
(a) Average correct-decision probability in the large-$\gamma$ regime for the asymmetric four-point constellation from Appendix A. The Monte Carlo estimate is compared with the angular limit predicted by Theorem 2, namely (76), and shown in the legend as “Angular limit (Thm. 2).” The simulated curve decreases steadily over the tested range and approaches the predicted plateau as the noise scale becomes large.
(b) Conditional correct-decision probabilities in the same large-$\gamma$ regime. The Monte Carlo estimates for $P_{1},P_{2},P_{3},$ and $P_{4}$ are shown explicitly in the legend, together with the corresponding angular (conditional) limits from Theorem 2 and Appendix A. The marker encoding is: black circles for $P_{1}=(0,0)$, blue squares for $P_{2}=(1,0)$, red upward triangles for $P_{3}=(0,1)$, and green downward triangles for $P_{4}=(0,-1)$. The numerical curves clearly approach the predicted limits in (75). In particular, the collapse of $P_{c}(P_{1};\gamma)$ toward zero provides direct numerical evidence for geometric collapse.
Figure 2: Finite-$\gamma$ validation of the large-noise theory in Theorem 2 for the asymmetric four-point constellation from Appendix A. Panel (a) shows that the average correct-decision probability approaches the angular limit $1/4$. Panel (b) shows that the corresponding symbol-level probabilities converge to their individual angular limits determined by the Voronoi recession cones. These results numerically confirm that, in the large-noise regime, reliability is governed by angular geometry rather than by bounded Euclidean distance information.
Fig. 1 summarizes the results. In Fig. 1(a), the Monte Carlo average symbol error probability remains strictly below the exact upper bound throughout the tested range, as expected. Moreover, the exact bound and its leading-order asymptotic approximation are nearly indistinguishable over the entire displayed small-$\gamma$ regime, indicating that the reciprocal-distance-spectrum law predicted by Theorem 1 emerges already at finite and practically observable noise levels. Although the upper bound is not numerically tight, it correctly captures the small-noise scaling and the relevant geometric dependence.
Fig. 1(b) further resolves the symbol-level behavior by plotting the conditional symbol error probabilities together with the corresponding exact conditional bounds and leading-order asymptotic lines. The ordering predicted in Appendix A is clearly observed:
$P_{e}(P_{1};\gamma)>P_{e}(P_{2};\gamma)>P_{e}(P_{3};\gamma)=P_{e}(P_{4};\gamma).$ (72)
Throughout the tested range, this numerical separation cannot be attributed to nearest-neighbor geometry alone, since all four points share the same minimum distance. Instead, it is consistent with the reciprocal-distance coefficients in Theorem 1, thereby illustrating that under isotropic Cauchy noise, even the high-reliability regime retains sensitivity to the global constellation geometry rather than only to $d_{\min}$.
Taken together, these simulations support two conclusions. First, Theorem 1 provides a finite-$\gamma$ upper bound with clear geometric content. Second, its leading-order reciprocal-distance descriptor already captures the correct qualitative reliability structure at the symbol level in the small-noise regime.
V-B Finite-$\gamma$ Validation of Theorem 2
We next validate the large-noise prediction in Theorem 2 using the same asymmetric four-point constellation from Appendix A,
$\mathcal{X}=\{(0,0),(1,0),(0,1),(0,-1)\}.$ (73)
For each
$\gamma\in\{$ 0.2, 0.3, 0.5, 0.7, 1, 1.5, $2,$03,05,07,1,15,2, (74)
$3,5,8,12,20,30\},$
we estimate the correct-decision probabilities by Monte Carlo simulation and compare them with the angular limits predicted by Theorem 2.
For this constellation, Appendix A gives the conditional large-noise limits
$\lim_{\gamma\to\infty}P_{c}(P_{1};\gamma)$ $=0,$ (75)
$\lim_{\gamma\to\infty}P_{c}(P_{2};\gamma)$ $=\frac{1}{4},$
$\lim_{\gamma\to\infty}P_{c}(P_{3};\gamma)=\lim_{\gamma\to\infty}P_{c}(P_{4};\gamma)$ $=\frac{3}{8},$
and hence the average correct-decision probability satisfies
$\lim_{\gamma\to\infty}P_{c}(\gamma)=\frac{1}{4}.$ (76)
Fig. 2 summarizes the numerical behavior. In Fig. 2(a), the Monte Carlo average correct-decision probability decreases steadily as $\gamma$ grows and approaches the predicted angular limit $1/4$. Although the convergence is naturally gradual, the trend is already clear on the finite-$\gamma$ range shown here, and the Monte Carlo values stabilize near the theorem-predicted plateau for the largest noise scales.
Fig. 2(b) further resolves the symbol-level behavior. The four conditional correct-decision probabilities approach distinct angular limits, exactly as predicted by Theorem 2 and Appendix A. In particular, the symbol $P_{1}=(0,0)$ becomes asymptotically unreliable, with its correct-decision probability tending to zero. This is the geometric-collapse phenomenon identified in Corollary 1. By contrast, $P_{2}$ approaches the nonzero limit $1/4$, while $P_{3}$ and $P_{4}$ approach the common limit $3/8$ and remain nearly indistinguishable throughout the simulation, consistent with the symmetry of the constellation.
These results provide finite-$\gamma$ numerical evidence for the large-noise geometric mechanism developed in Theorem 2. As the noise scale increases, bounded Euclidean offsets among constellation points lose influence, and the symbol-level reliability is governed instead by the angular measures of the corresponding Voronoi recession cones. Together with the small-noise validation in Section V-A, this experiment numerically supports the regime-dependent transition from distance-based to angle-based reliability descriptors under isotropic multivariate Cauchy noise.
VI Conclusion
This paper developed a geometric asymptotic perspective on one-shot detection under isotropic multivariate Cauchy noise. Although isotropic Cauchy noise induces the same Euclidean Voronoi decision regions as Gaussian noise, the corresponding reliability laws differ fundamentally because the probability mass assigned to those regions is redistributed by the heavy-tailed noise geometry. In the small-noise regime, we established a reciprocal distance-spectrum upper bound showing that error probabilities retain sensitivity to the full constellation geometry. In the large-noise regime, we proved that the correct-decision probability converges to an angular limit determined solely by the Voronoi recession cone. Together, these results formalize a regime-dependent transition from distance-based to angle-based reliability descriptors. The worked examples in Appendices A and B, together with the finite-$\gamma$ numerical validations in Section V, show that this transition can lead either to highly nonuniform symbol reliability, including complete geometric collapse, or to symmetric limiting behavior in standard constellations such as 4QAM.
Appendix A Geometric Collapse: A Four-Point Example
To concretely illustrate Theorems 1 and 2, we analyze the four-point constellation
$\mathcal{X}=\{P_{1},P_{2},P_{3},P_{4}\}\subset\mathbb{R}^{2},$ (77)
with
$P_{1}$ $=(0,0),$ $P_{2}$ $=(1,0),$ (78)
$P_{3}$ $=(0,1),$ $P_{4}$ $=(0,-1).$
This example is deliberately minimal: all points share the same nearest-neighbor distance
$d_{\min}=1,$ (79)
yet their large-noise behavior under isotropic Cauchy noise differs substantially because their Voronoi recession cones have different angular measures.
The AWGN column is included only as a qualitative comparison: under Gaussian noise, the nearest-neighbor structure largely governs the high-reliability regime, whereas under the isotropic Cauchy model the small-noise upper bound depends on the full reciprocal distance spectrum and the large-noise limit depends only on angular measures.
A-A Small-Noise Limit: Explicit Distance-Spectrum Evaluation
We first compute the pairwise distances:
$d_{12}=d_{13}=d_{14}$ $=1,$
$d_{23}=d_{24}$ $=\sqrt{2},$
$d_{34}$ $=2.$ (80)
By Theorem 1, for each transmitted symbol $P_{i}$,
$P_{e}(P_{i};\gamma)\leq\frac{2\gamma}{\pi}\sum_{j\neq i}\frac{1}{d_{ij}}+O(\gamma^{3}),\qquad\gamma\to 0.$ (81)
For $P_{1}=(0,0)$, all three competitors are at unit distance, so
$P_{e}(P_{1};\gamma)\leq\frac{2\gamma}{\pi}\left(1+1+1\right)+O(\gamma^{3})=\frac{6}{\pi}\gamma+O(\gamma^{3}).$ (82)
For $P_{2}=(1,0)$, the distances to $P_{1},P_{3},P_{4}$ are $1,\sqrt{2},\sqrt{2}$, respectively. Hence
$P_{e}(P_{2};\gamma)$ $\leq\frac{2\gamma}{\pi}\left(1+\frac{1}{\sqrt{2}}+\frac{1}{\sqrt{2}}\right)+O(\gamma^{3})$ (83)
$=\frac{2}{\pi}(1+\sqrt{2})\gamma+O(\gamma^{3}).$
For $P_{3}=(0,1)$, the distances to $P_{1},P_{2},P_{4}$ are $1,\sqrt{2},2$, respectively. Therefore
$P_{e}(P_{3};\gamma)$ $\leq\frac{2\gamma}{\pi}\left(1+\frac{1}{\sqrt{2}}+\frac{1}{2}\right)+O(\gamma^{3})$ (84)
$=\frac{1}{\pi}(3+\sqrt{2})\gamma+O(\gamma^{3}).$
By symmetry, the same expression holds for $P_{4}=(0,-1)$:
$P_{e}(P_{4};\gamma)\leq\frac{1}{\pi}(3+\sqrt{2})\gamma+O(\gamma^{3}).$ (85)
These calculations reproduce exactly the isotropic Cauchy small-$\gamma$ column in Fig. 3(b). In particular, although all four points share the same minimum distance, their reciprocal distance sums differ, so Theorem 1 already distinguishes them at leading order.
[…]
(a) Geometric properties
Constellation Point AWGN Isotropic Cauchy Isotropic Cauchy
$x$ (small $\sigma$) (small $\gamma$) ($\gamma\to\infty$, $P_{e}$)
$P_{1}=(0,0)$ $\approx 3\,Q\!\left(\frac{1}{2\sigma}\right)$ $\leq\frac{6}{\pi}\gamma+O(\gamma^{3})$ 11
$P_{2}=(1,0)$ $\approx Q\!\left(\frac{1}{2\sigma}\right)$ $\leq\frac{2}{\pi}(1+\sqrt{2})\gamma+O(\gamma^{3})$ $3/4$
$P_{3}=(0,1)$ $\approx Q\!\left(\frac{1}{2\sigma}\right)$ $\leq\frac{1}{\pi}(3+\sqrt{2})\gamma+O(\gamma^{3})$ $5/8$
$P_{4}=(0,-1)$ $\approx Q\!\left(\frac{1}{2\sigma}\right)$ $\leq\frac{1}{\pi}(3+\sqrt{2})\gamma+O(\gamma^{3})$ $5/8$
(b) Quantitative reliability descriptors
Figure 3: Geometric and quantitative analysis for the illustrative four-point constellation $\mathcal{X}$ (see Appendix A). (a) Left: Euclidean Voronoi cells. Right: recession cones and their angular patches on $\mathbb{S}^{1}$. The point $P_{1}$ has a degenerate cone of zero angular measure, whereas $P_{2},P_{3},P_{4}$ retain nontrivial patches with opening angles $\pi/2$, $3\pi/4$, and $3\pi/4$, respectively. (b) The corresponding theoretical reliability descriptors showcasing the quantitative transition from Gaussian minimum-distance behavior to Cauchy distance-spectrum and angular limits.
A-B Large-Noise Limit: Explicit Angular-Patch Evaluation
We now verify Theorem 2 by computing the recession cone
$\mathsf{K}(x)=\bigcap_{c\in\mathcal{X}\setminus\{x\}}\{u\in\mathbb{R}^{2}:\ u^{\mathsf{T}}(c-x)\leq 0\}$ (86)
for each constellation point. In two dimensions, Theorem 2 reduces to
$\lim_{\gamma\to\infty}P_{c}(x;\gamma)=\frac{\text{arc length of }A_{x}}{2\pi},\qquad A_{x}=\mathsf{K}(x)\cap\mathbb{S}^{1}.$ (87)
1) Point $P_{1}=(0,0)$
The competitor offsets are
$P_{2}-P_{1}$ $=(1,0),$ (88)
$P_{3}-P_{1}$ $=(0,1),$
$P_{4}-P_{1}$ $=(0,-1).$
Hence the cone constraints are
$u_{x}\leq 0,\quad u_{y}\leq 0,\quad-u_{y}\leq 0.$ (89)
Equivalently,
$u_{x}\leq 0,\quad u_{y}=0.$ (90)
Thus $\mathsf{K}(P_{1})$ is the negative $x$-axis ray, and the angular patch $A_{P_{1}}$ consists of a single point on $\mathbb{S}^{1}$. Its arc length is therefore zero, so
$\lim_{\gamma\to\infty}P_{c}(P_{1};\gamma)=0,\qquad\lim_{\gamma\to\infty}P_{e}(P_{1};\gamma)=1.$ (91)
This is the geometric-collapse phenomenon predicted by Corollary 1.
2) Point $P_{2}=(1,0)$
The competitor offsets are
$P_{1}-P_{2}$ $=(-1,0),$ (92)
$P_{3}-P_{2}$ $=(-1,1),$
$P_{4}-P_{2}$ $=(-1,-1).$
Hence
$u_{x}\leq 0,\quad-u_{x}+u_{y}\leq 0,\quad-u_{x}-u_{y}\leq 0,$ (93)
which is equivalent to
$u_{x}\geq 0,\quad|u_{y}|\leq u_{x}.$ (94)
Therefore $\mathsf{K}(P_{2})$ is a wedge centered on the positive $x$-axis with opening angle $\pi/2$, so its spherical patch has arc length $\pi/2$. By (87),
$\lim_{\gamma\to\infty}P_{c}(P_{2};\gamma)=\frac{\pi/2}{2\pi}=\frac{1}{4},\qquad\lim_{\gamma\to\infty}P_{e}(P_{2};\gamma)=\frac{3}{4}.$ (95)
3) Point $P_{3}=(0,1)$
The competitor offsets are
$P_{1}-P_{3}$ $=(0,-1),$ (96)
$P_{2}-P_{3}$ $=(1,-1),$
$P_{4}-P_{3}$ $=(0,-2).$
Hence
$u_{y}\leq 0,\quad u_{x}-u_{y}\leq 0,\quad-2u_{y}\leq 0,$ (97)
which reduces to
$u_{y}\geq 0,\quad u_{x}\leq u_{y}.$ (98)
Thus $\mathsf{K}(P_{3})$ is the wedge spanning angles from $\pi/4$ to $\pi$, whose opening angle is $3\pi/4$. Therefore
$\lim_{\gamma\to\infty}P_{c}(P_{3};\gamma)=\frac{3\pi/4}{2\pi}=\frac{3}{8},\qquad\lim_{\gamma\to\infty}P_{e}(P_{3};\gamma)=\frac{5}{8}.$ (99)
4) Point $P_{4}=(0,-1)$
By symmetry,
$P_{1}-P_{4}$ $=(0,1),$ (100)
$P_{2}-P_{4}$ $=(1,1),$
$P_{3}-P_{4}$ $=(0,2),$
so the cone constraints are
$u_{y}\leq 0,\quad u_{x}+u_{y}\leq 0.$ (101)
Equivalently,
$u_{y}\leq 0,\quad u_{x}\leq-u_{y}.$ (102)
Hence $\mathsf{K}(P_{4})$ is the reflected wedge spanning angles from $\pi$ to $7\pi/4$, again with opening angle $3\pi/4$. Thus
$\lim_{\gamma\to\infty}P_{c}(P_{4};\gamma)=\frac{3\pi/4}{2\pi}=\frac{3}{8},\qquad\lim_{\gamma\to\infty}P_{e}(P_{4};\gamma)=\frac{5}{8}.$ (103)
Under equiprobable signaling, symmetry implies that the average correct-decision probability also satisfies
$\lim_{\gamma\to\infty}P_{c}(\gamma)$ $=\frac{1}{4}\left(0+\frac{1}{4}+\frac{3}{8}+\frac{3}{8}\right)$ (104)
$=\frac{1}{4},$
and hence
$\lim_{\gamma\to\infty}P_{e}(\gamma)=1-\frac{1}{4}=\frac{3}{4}.$ (105)
This example shows explicitly that the small-noise and large-noise regimes are governed by fundamentally different geometric descriptors. In the high-reliability regime, the relevant quantity is the reciprocal distance sum appearing in Theorem 1; in the infinite-noise regime, all bounded distance information disappears and only the angular patch measure from Theorem 2 survives.
Appendix B A Standard 4QAM Sanity Check
As a complementary sanity check, we briefly examine the standard 44QAM constellation
$\mathcal{Q}=\{Q_{1},Q_{2},Q_{3},Q_{4}\}\subset\mathbb{R}^{2},$ (106)
with
$Q_{1}$ $=(1,1),$ $Q_{2}$ $=(-1,1),$ (107)
$Q_{3}$ $=(-1,-1),$ $Q_{4}$ $=(1,-1).$
Unlike the asymmetric four-point example in Appendix A, this constellation is fully symmetric. As a result, both the small-noise distance descriptor and the large-noise angular descriptor are identical for all four symbols.
B-A Small-Noise Regime
For any fixed point, say $Q_{1}=(1,1)$, the distances to the other three points are
$\|Q_{1}-Q_{2}\|=2,\quad\|Q_{1}-Q_{4}\|=2,\quad\|Q_{1}-Q_{3}\|=2\sqrt{2}.$ (108)
Hence, by Theorem 1,
$P_{e}(Q_{1};\gamma)$ $\leq\frac{2\gamma}{\pi}\left(\frac{1}{2}+\frac{1}{2}+\frac{1}{2\sqrt{2}}\right)+O(\gamma^{3})$ (109)
$=\frac{1}{\pi}\left(2+\frac{1}{\sqrt{2}}\right)\gamma+O(\gamma^{3}),\quad\gamma\to 0.$
By symmetry, the same bound holds for $Q_{2},Q_{3},Q_{4}$. Thus all symbols have the same leading-order small-noise behavior.
B-B Large-Noise Regime
We next evaluate the recession cones. For $Q_{1}=(1,1)$, the competitor offsets are
$Q_{2}-Q_{1}$ $=(-2,0),$ (110)
$Q_{3}-Q_{1}$ $=(-2,-2),$
$Q_{4}-Q_{1}$ $=(0,-2).$
Therefore, the cone constraints are
$2u_{x}\leq 0,\quad-2u_{x}-2u_{y}\leq 0,\quad-2u_{y}\leq 0,$ (111)
which implies
$u_{x}\geq 0,\quad u_{y}\geq 0.$ (112)
Hence $\mathsf{K}(Q_{1})$ is the first quadrant, whose angular patch on $\mathbb{S}^{1}$ has arc length $\pi/2$. By Theorem 2,
$\lim_{\gamma\to\infty}P_{c}(Q_{1};\gamma)=\frac{\pi/2}{2\pi}=\frac{1}{4},\quad\lim_{\gamma\to\infty}P_{e}(Q_{1};\gamma)=\frac{3}{4}.$ (113)
Again by symmetry, the same limit holds for all four symbols:
$\lim_{\gamma\to\infty}P_{c}(Q_{i};\gamma)=\frac{1}{4},\quad\lim_{\gamma\to\infty}P_{e}(Q_{i};\gamma)=\frac{3}{4},$ (114)
for $i=1,2,3,4$. Under equiprobable signaling, the average correct-decision probability therefore satisfies
$\lim_{\gamma\to\infty}P_{c}(\gamma)=\frac{1}{4},\quad\lim_{\gamma\to\infty}P_{e}(\gamma)=\frac{3}{4}.$ (115)
This example serves as a useful baseline: for a highly symmetric constellation such as 44QAM, both the reciprocal distance sum in Theorem 1 and the angular patch measure in Theorem 2 are identical across symbols. In contrast, the asymmetric example in Appendix A shows that these descriptors can vary substantially from point to point, leading to nonuniform reliability in both the small-noise and large-noise regimes.